# OmnimatteZero: Fast Training-free Omnimatte with Pre-trained Video Diffusion Models

### v2

@@ Abstract.
journal: TOGjournalvolume: 00journalnumber: 0article: 000publicationmonth: 8submissionid: papers_1753
@@ 2.3. Diffusion Transformers for Video Generation
[…]
[motivation figure]motivation figure
Figure 2. Comparison of object removal results using (a) vanilla image inpainting extended to video, (b) Vid2Vid zero-shot inpainting, and (c) our guidance-based approach. Vanilla methods fail to maintain temporal consistency or clean background reconstruction, while our method achieves coherent, artifact-free inpainting across frames.
[…]
[method figure]method figure
Figure 3. Overview of our Object Removal strategy in OmnimatteZero. (a) We first identify potential background correspondences across frames. (b) Temporal Attention Guidance (TAG): Temporal attention scores between a foreground point and its background correspondences are replaced with the average attention between all background pairs, promoting consistent inpainting across time. (c) Spatial Attention Guidance (SAG): Within a frame, the attention from a foreground point to nearby background points is adjusted to reflect the mean attention among background points themselves, improving inpainting quality when temporal context is unavailable.
@@ 5.1. Object Removal
Identifying Potential Background Patches from Neighboring Frames. To accurately inpaint masked areas in a specific frame, it is essential to identify corresponding background patches from neighboring frames, while preserving awareness of the video’s underlying 3D structure. This task can be effectively addressed using Track-Any-Point (TAP-Net) (Doersch et al., 2023), a real-time tracking framework. TAP-Net captures spatial and temporal coherence by tracking points from one frame to their corresponding locations across subsequent frames, as illustrated in Figure 3a, where points from frame $f_{i}$ are matched across other frames. Formally, given a point $p$ on the object in frame $i$, TAP-Net identifies its corresponding point set among frames $C=\{C_{j}|j\neq i\}$. To ensure that only background regions are considered, we discard points in $C$ that intersect the object mask $\mathcal{M}$ in their respective frames.
Temporal Attention Guidance (TAG): To effectively guide the denoising process toward inpainting using neighboring frames, we explicitly set the temporal attention score between a foreground point $p$ (in frame $i$) and each of its background correspondences $c\in C$ to be the mean temporal attention observed between all distinct pairs of corresponding background points (see Figure 3b). Formally, we first compute the mean temporal attention
@@
Since the attention is bidirectional, the denominator $|C|(|C|-1)$ correctly enumerates all ordered pairs $(c_{j},c_{k})$ where $j\neq k$, ensuring that each directional interaction is counted exactly once.
@@
Spatial Attention Guidance (SAG): To further improve spatial inpainting within each frame and to address cases where a point lacks correspondences in other frames, we reinforce spatial attention within the same frame. Specifically, for any foreground point $p$ in frame $i$, we set its spatial attention scores with surrounding background points $S$ to be the mean spatial attention computed among all pairs of background points within the same frame (see Figure 3c). Formally:
@@ 5.2. Removing associated object effects
(3) $A_{i}^{l}=\text{softmax}\left(\frac{Q_{i}^{l}(M_{obj}\odot K_{i}^{l})^{T}}{\sqrt{d}}\right)\in{\mathbb{R}}^{w\times h}$
where $Q^{l}_{i},K^{l}_{i}\in{\mathbb{R}}^{w\times h\times c}$ are the queries and keys respectively at layer $l$ and frame $i$, and $c$ is the number of channels of the latent.
[…]
[Description]Description
Figure 4. Self-attention maps from (a) LTX Video diffusion model and (b) Stable Diffusion (image based). The spatio-temporal video latent ”attends to object associated effects” (e.g., shadow, reflection) where, image models struggles to capture these associations.
We then compute a mask per-frame, $\mathcal{M}_{i}\in[0,1]^{w\times h\times c}$, which is aggregated across all diffusion layers:
@@
[…]
[Description]Description
Figure 5. (a) Foreground Extraction: The target object is extracted by latent code arithmetic, subtracting the background video encoding from the object+background latent (Latent Diff). This initially results in distortions, which are later corrected by replacing pixel values using the original video and a user-provided mask (Latent Diff + Pixel injection). (b) Layer Composition: The extracted object layer is added to a new background latent (Latent Addition). To improve blending, a few steps of noising-denoising are applied, yielding a more natural integration of the object into the new scene (Latent Addition + Refinement). See video examples in the supplemental material
@@
ObjectDrop (Winter et al., 2024) 28.05 0.124 34.22 0.083 31.14 0.104 - -
Video Repaint [LTXVideo]* (Lugmayr et al., 2022) 20.13 0.252 21.15 0.289 20.64 0.271 0 0.4
Video Repaint [Wan2.1]* (Lugmayr et al., 2022) 21.44 0.244 24.16 0.261 22.8 0.253 0 32
Temporal Enhance [LTXVideo]* (Luo et al., 2025) 21.33 0.248 23.01 0.281 22.8 0.265 0 0.04
Temporal Enhance [Wan2.1]* (Luo et al., 2025) 21.93 0.222 24.26 0.253 23.01 0.237 0 3.2
Lumiere inpainting (Bar-Tal et al., 2024) 26.62 0.148 31.46 0.157 29.04 0.153 - 9
Propainter (Zhou et al., 2023) 27.44 0.114 34.67 0.056 31.06 0.085 - 0.083
DiffuEraser (Zhou et al., 2023) 29.51 0.105 35.19 0.048 32.35 0.077 - 0.8
Ominmatte (Lu et al., 2021) 21.76 0.239 26.81 0.207 24.29 0.223 3 2.5
D2NeRF (Wu et al., 2024) - - 34.99 0.113 - - 3 2.2
LNA (Kasten et al., 2021) 23.10 0.129 - - - - 8.5 0.4
OmnimatteRF (Lin et al., 2023) 33.86 0.017 40.91 0.028 37.38 0.023 6 3.5
Generative Omnimatte (Lee et al., 2024) 32.69 0.030 44.07 0.010 38.38 0.020 - 9
OmnimatteZero [LTXVideo] (Ours) 35.11 0.014 44.97 0.010 40.04 0.012 0 0.04
OmnimatteZero [Wan2.1] (Ours) 34.12 0.017 45.55 0.006 39.84 0.011 0 3.2
@@ 5.3. Foreground extraction
This correction preserves the object’s fidelity while maintaining its associated effects, resulting in high-quality object layers (see Figure 5a Latent Diff + Pixel Injection).
@@ 6. Experiments
[…]
[Description]Description
Figure 6. Qualitative Results: Object removal and background reconstruction. The first row shows input video frames with object masks, while the second row presents the reconstructed backgrounds. Our approach effectively removes objects while preserving fine details, reflections, and textures, demonstrating robustness across diverse scenes. Notice the removal of the cat’s reflection in the mirror and water, the shadow of the dog and bicycle (with the rider), and the bending of the trampoline when removing the jumpers. See video examples in the supplemental material
@@ 6.1. Background layer reconstruction
Datasets and Metrics: We evaluate our method on two standard Omnimatte datasets: Movies (Lin et al., 2023) and Kubric (Wu et al., 2024). These datasets provide object masks for each frame and ground truth background layers for evaluation. All methods are assessed using PSNR and LPIPS metrics to measure the accuracy of background reconstruction, along with comparisons of training and runtime efficiency on a single A100 GPU.
@@
[…]
[Description]Description
Figure 7. Qualitative Comparison: Foreground Extraction. Foreground extraction comparison between OmnimatteZero and OmnimatteRF (Lin et al., 2023). Our method accurately captures both the object and its associated effects, such as shadows and reflections, in contrast to OmnimatteRF, missing or distorting shadows (row 1) and reflections (row 2).
Qualitative results: Figure 6 presents qualitative results of our object removal method across various scenes. The first row shows input video frames with masked objects, while the second row displays the reconstructed backgrounds. The videos include (1) a cat running toward a mirror, (2) a static cat looking left and right, (3) a dog running toward a man, (4) two people walking, and (5) multiple people jumping on a trampoline. Our method effectively removes objects like people and animals even when similar objects and effects appear in the video, while preserving fine details and textures. Notably, in the first two columns, OmnimatteZero eliminates the cat without leaving mirror or water reflections. The last column further demonstrates its ability to remove objects while maintaining scene integrity, even correcting the trampoline’s bending effect after removing the jumper. Figure 9 10 presents a qualitative comparison of OmnimatteZero with SoTA omnimatte and video inpainting methods. Our method consistently produces cleaner and more temporally coherent background reconstructions with fewer artifacts and distortions, especially in challenging scenes involving motion, shadows, and texture continuity. The figures includes diverse videos such as dogs running on grass, a person walking past a building, a woman walking on the beach, swans in a lake and more. In the first three rows of Figure 10, OmnimatteZero reconstructs textured and complex backgrounds (e.g., grass, water, sand) without the distortions and ghosting present in OmnimatteRF and DiffuEraser (see red boxes). In all rows of Figure 9, competing methods leave visible shadow remnants, semantic drift, or object traces, while OmnimatteZero removes foregrounds cleanly and fills in semantically appropriate content.
[…]
[Description]Description
Figure 8. Qualitative Comparison: Layer Composition. The extracted foreground objects, along with their shadows and reflections, are seamlessly integrated into diverse backgrounds, demonstrating the effectiveness of our approach in preserving visual coherence and realism across different scenes. See video examples in the supplemental material
@@ 6.3. Layer composition
The objective here is to take the extracted foreground layers and seamlessly composite them onto new background layers. Figures 8 and 11 presents the results of OmnimatteZero, highlighting its ability to maintain object integrity, shadows, and reflections while enabling smooth re-composition across different scenes. Notably, it demonstrates how the model adaptively integrates the inserted object into the environment, such as enhancing the cat’s reflection in the clear water stain or adjusting the man’s appearance to match the night lighting.
6.4. Acknowledgement
@@
Appendix A Implementation Details:
@@ B.3. Impact of Effect-Aware Masking
[…]
[Description]Description
@@
Figure 9. Qualitative comparison with state-of-the-art video inpainting methods. OmnimatteZero produces clean, temporally consistent background reconstructions across a diverse range of scenes, including dynamic motion, shadows, reflections, and fine textures. Compared to other methods, it avoids common artifacts such as ghosting, blur, and shadow remnants (highlighted in red boxes), successfully filling in complex backgrounds across varying temporal and spatial contexts. Gray frames indicate failure cases where the inpainting method returned an error. See the supplementary material for full videos.
[…]
[Description]Description
@@
Figure 10. Qualitative comparison with state-of-the-art video inpainting methods. OmnimatteZero achieves clean, temporally consistent background reconstructions across diverse scenes, preserving fine textures, shadows, and reflections. Compared to prior methods, it avoids common artifacts such as ghosting, blur, and residual shadows (highlighted in red). Gray frames indicate failure cases where the method returned an error. See supplementary material for full video results.
[…]
[Description]Description
@@
Figure 11. Qualitative Comparison: Layered Composition. From left to right: the original video frame, the extracted foreground (including shadows and reflections), the target background, and our final composite. Across four diverse examples (a swan, a cat, a bicyclist, and a dog), our method preserves accurate shadows, reflections, and object–scene coherence, yielding visually seamless results. See the supplementary material for full videos.

### v3

@@ Abstract.
$c$s: Computing methodologiesccs: Computing methodologies Machine learning algorithmsjournalyear: 2025copyright: ccconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763917isbn: 979-8-4007-2137-3/2025/12submissionid: papers_1753
@@ 2.3. Diffusion Transformers for Video Generation
Figure 2. Comparison of object removal results using (a) vanilla image inpainting extended to video, (b) Vid2Vid zero-shot inpainting, and (c) our guidance-based approach. Vanilla methods fail to maintain temporal consistency or clean background reconstruction, while our method achieves coherent, artifact-free inpainting across frames.: motivation figure
Figure 3. Overview of our Object Removal strategy in OmnimatteZero. (a) We first identify potential background correspondences across frames. (b) Temporal Attention Guidance (TAG): Temporal attention scores between a foreground point and its background correspondences are replaced with the average attention between all background pairs, promoting consistent inpainting across time. (c) Spatial Attention Guidance (SAG): Within a frame, the attention from a foreground point to nearby background points is adjusted to reflect the mean attention among background points themselves, improving inpainting quality when temporal context is unavailable.: method figure
@@ 5. Method: OmnimatteZero
In summary, our pipeline expands masks to capture object effects, inpaints them for clean backgrounds, isolates foregrounds via latent arithmetic, and seamlessly recomposes them onto new scenes.
@@ 5.1. Object Removal
Identifying Potential Background Patches from Neighboring Frames
To accurately inpaint masked areas in a specific frame, it is essential to identify corresponding background patches from neighboring frames, while preserving awareness of the video’s underlying 3D structure. This task can be effectively addressed using Track-Any-Point (TAP-Net) (Doersch et al., 2023), a real-time tracking framework. TAP-Net captures spatial and temporal coherence by tracking points from one frame to their corresponding locations across subsequent frames, as illustrated in Figure 3a, where points from frame $f_{i}$ are matched across other frames. Formally, given a point $p$ on the object in frame $i$, TAP-Net identifies its corresponding point set among frames $C=\{C_{j}|j\neq i\}$. To ensure that only background regions are considered, we discard points in $C$ that intersect the object mask $\mathcal{M}$ in their respective frames.
Temporal Attention Guidance (TAG)
To effectively guide the denoising process toward inpainting using neighboring frames, we explicitly set the temporal attention score between a foreground point $p$ (in frame $i$) and each of its background correspondences $c\in C$ to be the mean temporal attention observed between all distinct pairs of corresponding background points (see Figure 3b). Formally, we first compute the mean temporal attention
@@
Since the attention is bidirectional, the denominator |C|(|C|1)|C|(|C|-1) correctly enumerates all ordered pairs $(c_{j},c_{k})$ where $j\neq k$, ensuring that each directional interaction is counted exactly once.
@@
Spatial Attention Guidance (SAG)
To further improve spatial inpainting within each frame and to address cases where a point lacks correspondences in other frames, we reinforce spatial attention within the same frame. Specifically, for any foreground point $p$ in frame $i$, we set its spatial attention scores with surrounding background points $S$ to be the mean spatial attention computed among all pairs of background points within the same frame (see Figure 3c). Formally:
@@
Figure 4. Self-attention maps from (a) LTX Video diffusion model and (b) Stable Diffusion (image based). The spatio-temporal video latent ”attends to object associated effects” (e.g., shadow, reflection) where, image models struggles to capture these associations.: Description
@@ 5.2. Removing Associated Object Effects
(3) $A_{i}^{l}=\text{softmax}\left(\frac{Q_{i}^{l}(M_{obj}\odot K_{i}^{l})^{T}}{\sqrt{c}}\right)\in{\mathbb{R}}^{N\times N}$
where $Q^{l}_{i},K^{l}_{i}\in{\mathbb{R}}^{w\times h\times c}$ are the queries and keys respectively at layer $l$ and frame $i$, $N=w\times h$ and $c$ is the number of channels of the latent.
We then compute a soft-mask per-frame, $\mathcal{M}_{i}\in[0,1]^{w\times h\times c}$, which is aggregated across all diffusion layers:
@@
Figure 5. (a) Foreground Extraction: The target object is extracted by latent code arithmetic, subtracting the background video encoding from the object+background latent (Latent Diff). This initially results in distortions, which are later corrected by replacing pixel values using the original video and a user-provided mask (Latent Diff + Pixel injection). (b) Layer Composition: The extracted object layer is added to a new background latent (Latent Addition). To improve blending, a few steps of noising-denoising are applied, yielding a more natural integration of the object into the new scene (Latent Addition + Refinement). See video examples in the supp material.: Description
@@
ObjectDrop (Winter et al., 2024) 28.05 0.124 - 34.22 0.083 - 31.14 0.104 - - -
Video Repaint [LTXVideo]* (Lugmayr et al., 2022) 20.13 0.252 - 21.15 0.289 - 20.64 0.271 - 0 0.4
Video Repaint [Wan2.1]* (Lugmayr et al., 2022) 21.44 0.244 - 24.16 0.261 - 22.8 0.253 - 0 32
Temporal Enhance [LTXVideo]* (Luo et al., 2025) 21.33 0.248 - 23.01 0.281 - 22.8 0.265 - 0 0.04
Temporal Enhance [Wan2.1]* (Luo et al., 2025) 21.93 0.222 - 24.26 0.253 - 23.01 0.237 - 0 3.2
Lumiere inpainting (Bar-Tal et al., 2024) 26.62 0.148 - 31.46 0.157 - 29.04 0.153 - - 9
Propainter (Zhou et al., 2023) 27.44 0.114 - 34.67 0.056 - 31.06 0.085 - - 0.083
DiffuEraser (Zhou et al., 2023) 29.51 0.105 - 35.19 0.048 - 32.35 0.077 - - 0.8
Ominmatte (Lu et al., 2021) 21.76 0.239 0.736 26.81 0.207 0.831 24.29 0.223 0.783 3 2.5
D2NeRF (Wu et al., 2024) - - - 34.99 0.113 0.887 - - - 3 2.2
LNA (Kasten et al., 2021) 23.10 0.129 0.847 - - - - - - 8.5 0.4
OmnimatteRF (Lin et al., 2023) 33.86 0.017 0.981 40.91 0.028 0.970 37.38 0.023 0.975 6 3.5
Generative Omnimatte (Lee et al., 2024) 32.69 0.030 0.989 44.07 0.010 0.981 38.38 0.020 0.985 - 9
OmnimatteZero [LTXVideo] (Ours) 35.11 0.014 0.992 44.97 0.010 0.988 40.04 0.012 0.990 0 0.04
OmnimatteZero [Wan2.1] (Ours) 34.12 0.017 0.993 45.55 0.006 0.987 39.84 0.011 0.990 0 3.2
@@ 5.3. Foreground Extraction
This correction preserves the object’s fidelity while maintaining its associated effects, resulting in high-quality object layers (see Figure 5a Latent Diff + Pixel Injection). A comparison of OmnimatteZero with other baselines for foreground extraction is provided in the Supplemental.
Alpha-Channel Extraction: The alpha-channel is obtained from the soft mask $M_{i}$ (Eq. 4) by decoding it with the video VAE using only the spatial and temporal upsampling layers, skipping feature re-encoding. This ensures (1) softness, the output remains a soft mask that captures graded transitions around boundaries and effects, and (2) alignment, the decoded alpha-channel is upsampled to the original resolution, maintaining pixel-wise correspondence with the RGB frames. The result is a temporally and spatially consistent matte that preserves soft boundaries, shadows, and reflections, enabling realistic compositing and evaluation.
@@ 6. Experiments
Figure 6. Qualitative Results: Object removal and background reconstruction. The first row shows input video frames with object masks, while the second row presents the reconstructed backgrounds. Our approach effectively removes objects while preserving fine details, reflections, and textures, demonstrating robustness across diverse scenes. Notice the removal of the cat’s reflection in the mirror and water, the shadow of the dog and bicycle (with the rider), and the bending of the trampoline when removing the jumpers. See video examples in the supplemental material : Description
@@ 6.1. Background Layer Reconstruction
Datasets and Metrics: We evaluate our method on two standard Omnimatte datasets: Movies (Lin et al., 2023) and Kubric (Wu et al., 2024). These datasets provide object masks for each frame and ground truth background layers for evaluation. All methods are assessed using PSNR, SSIM, and LPIPS metrics to measure the accuracy of background reconstruction, along with comparisons of training and runtime efficiency on a single A100 GPU.
@@
Figure 7. Qualitative Comparison: Foreground Extraction. Foreground extraction comparison between OmnimatteZero and OmnimatteRF (Lin et al., 2023). Our method accurately captures both the object and its associated effects, such as shadows and reflections, in contrast to OmnimatteRF, missing or distorting shadows (row 1) and reflections (row 2). : Description
Qualitative results: Figure 6 presents qualitative results of our object removal method across various scenes. The first row shows input video frames with masked objects, while the second row displays the reconstructed backgrounds. The videos include (1) a cat running toward a mirror, (2) a static cat looking left and right, (3) a dog running toward a man, (4) two people walking, and (5) multiple people jumping on a trampoline. Our method effectively removes objects like people and animals even when similar objects and effects appear in the video, while preserving fine details and textures. Notably, in the first two columns, OmnimatteZero eliminates the cat without leaving mirror or water reflections. The last column further demonstrates its ability to remove objects while maintaining scene integrity, even correcting the trampoline’s bending effect after removing the jumper. Figure 10 11 presents a qualitative comparison of OmnimatteZero with SoTA omnimatte and video inpainting methods. Our method consistently produces cleaner and more temporally coherent background reconstructions with fewer artifacts and distortions, especially in challenging scenes involving motion, shadows, and texture continuity. The figures includes diverse videos such as dogs running on grass, a person walking past a building, a woman walking on the beach, swans in a lake and more. In the first three rows of Figure 11, OmnimatteZero reconstructs textured and complex backgrounds (e.g., grass, water, sand) without the distortions and ghosting present in OmnimatteRF and DiffuEraser (see red boxes). In all rows of Figure 10, competing methods leave visible shadow remnants, semantic drift, or object traces, while OmnimatteZero removes foregrounds cleanly and fills in semantically appropriate content.
Figure 8. Qualitative Comparison: Layer Composition. The extracted foreground objects, along with their shadows and reflections, are seamlessly integrated into diverse backgrounds, demonstrating the effectiveness of our approach in preserving visual coherence and realism across different scenes. See video examples in the supplemental material: Description
@@ 6.3. Layer Composition
The objective here is to take the extracted foreground layers and seamlessly composite them onto new background layers. Figures 8 and 12 presents the results of OmnimatteZero, highlighting its ability to maintain object integrity, shadows, and reflections while enabling smooth re-composition across different scenes. Notably, it demonstrates how the model adaptively integrates the inserted object into the environment, such as enhancing the cat’s reflection in the clear water stain or adjusting the man’s appearance to match the night lighting.
7. Conclusion and Limitations
In this work, we introduced OmnimatteZero, the first training-free approach for video omnimatte. By leveraging pre-trained video diffusion models together with spatial and temporal attention guidance, OmnimatteZero enables real-time object removal, foreground extraction, and seamless layer recomposition without any additional training or test-time optimization. Extensive experiments demonstrate that our method outperforms supervised and self-supervised baselines in both background reconstruction and multi-object handling, while achieving state-of-the-art runtime efficiency.
Nevertheless, several limitations remain. First, since our pipeline relies on VAE compression and decompression, the reconstructed video may exhibit slight deviations from the original input. Second, the quality of temporal correspondence depends on the accuracy of TAP, which can degrade under heavy occlusions or in low-resolution videos. Third, the fidelity of inpainting is ultimately bounded by the capabilities of the underlying video diffusion model, which may struggle in highly complex or unfamiliar scenes. We believe these limitations will be progressively mitigated as future video generative models continue to improve in visual fidelity, temporal consistency, and robustness. OmnimatteZero provides a fast, flexible, and practical framework that can readily incorporate such advances, paving the way for scalable, real-time video editing applications.
Acknowledgements.
@@
Appendix A Implementation Details
@@ B.3. Impact of Effect-Aware Masking
Figure 9. Qualitative examples of multi-object extraction. OmnimatteZero successfully separates multiple dynamic foreground objects (e.g., person, dog, puppies) along with their associated effects such as shadows. The results demonstrate that our method scales to complex scenes with several independently moving objects, preserving fine details and temporal consistency.: Description
@@
Appendix C Multi-Object Extraction
Figure 9 provides qualitative examples of multiple foreground extractions from a single video. It shows that OmnimatteZero can reliably isolate several dynamic objects together with their associated effects, such as shadows and reflections. Our method maintains temporal consistency and preserves fine details across all layers. This demonstrates the scalability of our framework to complex multi-object scenes. Full video results are provided in the supplementary material.
Appendix D Handling Object Occlusion
We further evaluated the ability of our method to handle object occlusion, a particularly challenging scenario for video decomposition. To this end, we constructed a benchmark of 20 videos containing significant occlusions between objects. In these experiments, our method successfully produced coherent object removal and background reconstruction in 60% of the cases. This substantially outperforms Generative-Omnimatte (Lee et al., 2024), which succeeded in only 30% of cases, and OmnimatteRF (Lin et al., 2023), which achieved just 1%. These results indicate that the proposed spatial and temporal attention guidance enables improved robustness to occlusions compared to prior approaches. Nonetheless, occlusion remains a difficult challenge, as heavy overlaps can degrade tracking accuracy and limit the available background context. We believe that future advancements in video generative models and more robust tracking mechanisms will further enhance performance in these scenarios.
Appendix E Comparison to Generative-Omnimatte
Here we provide further discussion and comparison with Generative-Omnimatte (Lee et al., 2024). Generative-Omnimatte is a supervised video diffusion method for object removal. While effective in certain cases, it requires annotated training data, per-video test-time optimization, and incurs several seconds of computation per frame. Moreover, it often struggles in scenes containing multiple objects and complex interactions. In contrast, our approach is training-free, requires no test-time optimization, and achieves real-time performance (0.04 seconds per frame on an A100 GPU). Thanks to the proposed spatial and temporal attention guidance, our method can robustly handle scenes with multiple objects while maintaining both temporal consistency and accurate background reconstruction. To further evaluate this difference, we conducted an additional comparison on 20 challenging videos, each containing more than five objects. In these settings, we attempted to remove several objects simultaneously. Our method succeeded in all cases, whereas Generative-Omnimatte failed on 90% of the videos, leaving visible artifacts or incomplete removals. These results highlight the practical advantage of our approach in complex, multi-object scenarios, where supervised and optimization-based methods face severe limitations.
@@ Appendix E Comparison to Generative-Omnimatte
Figure 10. Qualitative comparison with state-of-the-art video inpainting methods. OmnimatteZero produces clean, temporally consistent background reconstructions across a diverse range of scenes, including dynamic motion, shadows, reflections, and fine textures. Compared to other methods, it avoids common artifacts such as ghosting, blur, and shadow remnants (highlighted in red boxes), successfully filling in complex backgrounds across varying temporal and spatial contexts. Gray frames indicate failure cases where the inpainting method returned an error. See the supplementary material for full videos.: Description
@@
Figure 11. Qualitative comparison with state-of-the-art video inpainting methods. OmnimatteZero achieves clean, temporally consistent background reconstructions across diverse scenes, preserving fine textures, shadows, and reflections. Compared to prior methods, it avoids common artifacts such as ghosting, blur, and residual shadows (highlighted in red). Gray frames indicate failure cases where the method returned an error. See supplementary material for full video results.: Description
@@
Figure 12. Qualitative Comparison: Layered Composition. From left to right: the original video frame, the extracted foreground (including shadows and reflections), the target background, and our final composite. Across four diverse examples (a swan, a cat, a bicyclist, and a dog), our method preserves accurate shadows, reflections, and object–scene coherence, yielding visually seamless results. See the supplementary material for full videos. : Description